a nut store normally sells cashews $9 per pound and almonds for $3.50 per pound. But at the end of the month and the almonds had not sold well, so in order to sell 60 pounds of almonds, the manager decided to mix the 60 pounds of almonds with some cashews and sell the mixture for $7.50 per pound. How many pounds of cashews should be mixed with the almonds to ensure no change in the profit?
Answer by stanbon(75887)

a nut store normally sells cashews $9 per pound and almonds for $3.50 per pound. But at the end of the month and the almonds had not sold well, so in order to sell 60 pounds of almonds, the manager decided to mix the 60 pounds of almonds with some cashews and sell the mixture for $7.50 per pound. How many pounds of cashews should be mixed with the almonds to ensure no change in the profit?
----------------
Equations:
Quantity Eq: a + c = 60
Value Eq:::3.5a + 9c = 7.5*60
-----
Multiply thru 1st by 35
Multiply thru 2nd by 10
-------------------
35a + 35c = 35*60
35a + 90c = 75*60
-----
Subtract and solve for "c":
55c = 40*60
c = 43.64 lbs (amt of cashews needed in the mixture)
================
Cheers,
Stan H.
==============